Please help me with this question...
Pravesh and Solly usually go to a coffee shop during luch break and toss a coin to see who will pay. what is the probability that Pravesh will pay 3 days in a row?

Answer by stanbon(75887)

Pravesh and Solly usually go to a coffee shop during luch break and toss a coin to see who will pay. what is the probability that Pravesh will pay 3 days in a row?
----
Binomial Problem with n = 3 and P(pay) = 1/2
---
P(x = 3) = (1/2)^3 = 1/8
============================
Cheers,
Stan H.